OpenStudy (anonymous):

FAN AND MEDEAL Find the product.Simplify 1. 1/10 x 3/4 A 15/2 B 2/15 C 3/40 D 40/3 2. 4/5 OF 40 A 32 B 1/32 C 12 D 44/5 3. 1/3 OF 3 A 1/18 B 1/3 C 1 D 1/9 4. Betty had 1/7 yard of ribbon. She used 1/6 of it. How many yards of ribbon did she use? 5. Suppose you have 15 muffins. How many muffins are left after you give a friend 1/3 of them? A 12 muffins B 10 muffins C 45 muffins D 5 muffins

Nnesha (nnesha):

http://openstudy.com/study#/updates/54bc4095e4b0c3c33929ea5a if you didn't understand ask question please :)

myininaya (myininaya):

why did you close the previous thread?

myininaya (myininaya):

I want you to try to multiply the fractions by using the rule I posted for multiplying fractions... that is what is product of the tops? and what is the product of the bottoms?

myininaya (myininaya):

\[\frac{a}{b} \times \frac{c}{d}=\frac{a \cdot c}{b \cdot d }\] all you have to do according to this is multiply your top and multiply your bottoms tops stay on top bottoms stay on bottom